OpenStudy (anonymous):

Simplifying rational expressions. Simplify (2√5+3√7)^2 Simplify 4√6/√30 by rationalizing the denominator.

OpenStudy (anonymous):

\[(2\sqrt5+3\sqrt7)(2\sqrt5+3\sqrt7)\] is a start

OpenStudy (anonymous):

for the second one start with \[\frac{4}{\sqrt{5}}\] and then multiply by 1 in the form of \[\frac{4}{\sqrt{5}}\times \frac{\sqrt{5}}{\sqrt{5}}\]

OpenStudy (anonymous):

@satellite73 is the answer for the second one 4√5/5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i really dont know the first one though can you explain a little more?

OpenStudy (anonymous):

you have to multiply out using the distributive property \[(a+b)^2=(a+b)(a+b)=a^2+2ab+b^2\] or in this case it would be easier to compute at \[(a+b)^2=a^2+b^2+2ab\]

OpenStudy (anonymous):

\[(2\sqrt5+3\sqrt7)(2\sqrt5+3\sqrt7)\] \[=(2\sqrt{2})^2+(3\sqrt{7})^2+2\times 2\sqrt{5}\times 3\sqrt{7}\] is the first step

OpenStudy (anonymous):

oops typo there, should be \[=(2\sqrt{5})^2+(3\sqrt{7})^2+2\times 2\sqrt{5}\times 3\sqrt{7}\]

OpenStudy (anonymous):

then \[4\times 5+9\times 7+12\sqrt{5\times 7}\]

OpenStudy (anonymous):

i will let you finish it from there

OpenStudy (anonymous):

95√35 i think

OpenStudy (anonymous):

no

OpenStudy (anonymous):

If that's not it idk the answer.